Find work and internal energy for 1 mole of water if enthalpy of vaporization is 42.5 kjmol-1 at 100°c and 1atm. Neglect volume of liquid water and assume ideal behavior of vapor water

1
Expert's answer
2021-12-20T06:07:01-0500

H = U + ng . R . T

42.5 × 1000 = U + 1 × 8.314 × 373

42500 = U + 3101.122

U = ( 42500 - 3101.122 ) j / mol.

U= 39398.878 j / mol .

U = 39.398 kj / mol.

Where H is enthalpy of vapourisation

U = internal energy change .


Work done = ng . R .T .

= 1 × 8.314 × 373 .

= 3101.122 j

= 3.101122 kj .